OpenStudy (donnie1999):

What are the x-intercepts of the parabola x2 + 4x −12?

OpenStudy (donnie1999):

@H3LPN33DED

OpenStudy (donnie1999):

@aloud @LegendarySadist

OpenStudy (anonymous):

Factor x2+4x−12 x2+4x−12 =(x−2)(x+6) Answer: (x−2)(x+6)

OpenStudy (donnie1999):

these r the choices a. (-6,0), (2,0) b. (0,-6),(0,2) c. (0,6),(0,-2) d. (6,0),(-2,0)

OpenStudy (donnie1999):

@aloud

OpenStudy (epicteatime):

did you graph it?

OpenStudy (epicteatime):

it'll be easier if you do :p

OpenStudy (epicteatime):

:3

OpenStudy (donnie1999):

@Epicteatime Im unable to graph it

OpenStudy (epicteatime):

ah

OpenStudy (anonymous):

Factor x2+4x−12 x2+4x−12 =(x−2)(x+6) Answer: (x−2)(x+6)

OpenStudy (epicteatime):

well, try using wolfram. It helps :p Or download geogebra

OpenStudy (epicteatime):

just type the equation in :D

OpenStudy (epicteatime):

unless this is a multiple question

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

Its really simple, just look at where it intersects. So I graphed it

OpenStudy (epicteatime):

x intercepts are -6 and 2

OpenStudy (donnie1999):

im confused witch answer would it be a. (-6,0), (2,0) b. (0,-6),(0,2) c. (0,6),(0,-2) d. (6,0),(-2,0)

OpenStudy (anonymous):

hey @mathmate

OpenStudy (epicteatime):

I told you the answer xD

OpenStudy (epicteatime):

and showed you the graph :p YOU SHOULD BE SET BUD BUD :D

OpenStudy (anonymous):

d

OpenStudy (epicteatime):

wha, I thought it was A

OpenStudy (epicteatime):

donnie's taking his sweet time xD

OpenStudy (anonymous):

nope its D look how i solved it

OpenStudy (epicteatime):

:p welp

OpenStudy (epicteatime):

xD

OpenStudy (donnie1999):

lol im so confused is it a or d

OpenStudy (epicteatime):

use his answer donnie, I didn't use the equation xD

OpenStudy (anonymous):

look (x−2)(x+6) -2 and +6 so it would be D

OpenStudy (epicteatime):

Its a xD You have to break it down, and then find the x's

OpenStudy (epicteatime):

so aloud would be correct

OpenStudy (epicteatime):

if you're still confused, then we can help if you want xD

OpenStudy (mathmate):

"look (x−2)(x+6) -2 and +6 so it would be D" (ref. aloud) @Epicteatime @donnie1999 The product theorem says that if a*b=0, then a=0, or b=0, or a=0 and b=0. @aloud has correctly factored the equation as (x−2)(x+6)=0. So according to the product theorem, (x-2)=0 or (x+6)=0 Solving these equations would give x=-6, or x=2.

OpenStudy (epicteatime):

I told you, you're correct xD

OpenStudy (epicteatime):

oh wait, your a different guy, but yeah, same thing, your correct bud bud gg :p